Problem 2-22 (Algo) CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure [LO2-1, LO2-3, LO2-4, LO2-5, LO2-6]

Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (12,700 units × $20 per unit) $ 254,000
Variable expenses 127,000
Contribution margin 127,000
Fixed expenses 142,000
Net operating loss $ (15,000 )

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $57,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,700 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assu
Answer #1
Req 5A :
Current unit variable cost = Variable expenses / Units sold = 127000 / 12700 10
Revised unit variable cost = 10 - 3 7
Revised fixed expenses = 142000 + 57000 199000
Contribution margin per unit = Selling price - Unit variable cost = 20 - 7 13
CM ratio = Contribution margin per unit / Selling price = 13 / 20 65%
Break-even point in unit sales = Fixed costs / Contribution margin per unit = 199000 / 13 15308
Break even point in dollar sales = Fixed costs / CM ratio = 199000 / 65% 306154
Req 5B :
Not automated Automated
Total Per unit % Total Per unit %
Sales 400000 20 100% 400000 20 100%
Variable expenses 200000 10 50% 140000 7 35%
Contribution margin 200000 10 50% 260000 13 65%
Fixed expenses 142000 199000
Net operating income 58000 61000
Req 5C :
Answer : Yes
Explanation : If the company sells 20700 units, then the company will earn more net operating income if it automates its operations.